Paul Havlik promised his grandson Jamie that he would give him $6,300 7 years from today for graduating from high school. Assume money is worth 8% interest compounded semiannually.
  
What is the present value of this $6,300? (Do not round intermediate calculations. Round your answer to the nearest cent.)
Answer #1

We use the formula:
A=P(1+r/2)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

6,300=P*(1+0.08/2)^(2*7)

P=6,300/(1+0.08/2)^(2*7)

=6,300*0.577475082

=$3638.09(Approx).